OpenStudy (anonymous):

Change the given equation into slope-intercept form. 4x – 3y = -12

OpenStudy (anonymous):

Do you know the equation for the slope-intercept form?

OpenStudy (anonymous):

no this is a pretest so i havent learned it exactly yet

OpenStudy (anonymous):

Ok!

OpenStudy (anonymous):

The slope-intercept form is: y= ax+b

OpenStudy (anonymous):

As you can see, we're trying to isolate the y on one side and bring everything else to the other.

OpenStudy (anonymous):

The y is also positive and has only one of them ( So there is not 2y = ax + b for example)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

Let's do it step by step. Can you make your y by itself? (Bring the x over)

OpenStudy (anonymous):

ok. and no, i dont know how to do any of it. Im new to this kind of stuff and its confusing lol

OpenStudy (anonymous):

I'll give you a simple equation x + 1 = 5

OpenStudy (anonymous):

Would it be mathematically correct to say that x = 4? Because x+1-1 = 5-1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Exactly! So it is the same thing here but we add something harder. We have to move variables to the other side as well!

OpenStudy (anonymous):

4x -3y = 12 Would it be correct to say that 4x-4x-3y = 12-4x ???

OpenStudy (anonymous):

So -3y = 12-4x?

OpenStudy (anonymous):

ok but whats next?

OpenStudy (anonymous):

-3/2?

OpenStudy (anonymous):

Well as you can see, we've already isolated the y. What we want to do is get only one y and to make it positive.

OpenStudy (anonymous):

How do we get rid of the -3 in front of y?

OpenStudy (anonymous):

divide it by 3?

OpenStudy (anonymous):

By -3, close ;)

OpenStudy (anonymous):

So now... y = ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Well there is one y but I mean what is on the other side of y

OpenStudy (anonymous):

what is 12/-3 and -4x/-3

OpenStudy (anonymous):

-4 is the first one, the other us a decimal 1.33 repeating

OpenStudy (anonymous):

Correct. You could also leave it as a fraction.

OpenStudy (anonymous):

So y = 4/3x -4

OpenStudy (anonymous):

That would be your answer. 4/3 is the slope and -4 is the y-intercept

OpenStudy (pottersheep):

It has to be in y = mx + b form

OpenStudy (anonymous):

so it would be D?

OpenStudy (anonymous):

We just did it together....y = 4/3x -4

OpenStudy (anonymous):

You said it yourself.... -4!!!

OpenStudy (anonymous):

lol. sorry i meant C. thankyouuuu